OpenStudy (anonymous):

A marble is drawn from a bag containing 3 red marbles and 2 blue marbles and a number cube with faces numbered 1 through 6 is tossed. What is the probability that the marble drawn is blue and the number cube shows a number less than 3? A. 11/15 B. 1/4 C. 1/6 D. 2/15

OpenStudy (anonymous):

I think C

OpenStudy (anonymous):

explain? @mboga1

OpenStudy (anonymous):

It's B

OpenStudy (anonymous):

explain? @mboga2

OpenStudy (anonymous):

To find the probability of two events we need to multiply the probability of each event together. To do this we need to two individual probabilities (i.e. One for the bag and one for the cube.) Let's do the bag first: There are 5 total marbles and 2 "correct" blue marbles. To find the probabiltiy we do \[P=\frac{correct}{total}\] in this case that gives: \[P(blue)=\frac{2}{5}\] Similary for the cube there are 6 sides and only 2 "correct" number, so" \[P(<3)=\frac{2}{6}\] To get the total probability we do:\[P(AandB)=P(A) \times P(B)\]\[P(blue and<3)=P(blue) \times P(<3)\]\[P=\frac{2}{5}\times\frac{2}{6}=\frac{4}{30}=\frac{2}{15}\] (D)

OpenStudy (anonymous):

Thank you so much @twvogels !!!!